# Versatile Multimodal Controls for Whole-Body Talking Human Animation

### v2

@@ 3.3 Generating Photorealistic Talking and Moving Humans with Audio and Motion Control
3.3.1 Co-Speech Human Animation with Video Diffusion Model
@@
We extend an off-the-shelf human animation framework (Zhang et al., 2024b) by incorporating additional conditional signal, including audio, emotion labels, and blinking ratios. This enhancement enables the simultaneous generation of lip movements and whole-body animations.
@@
As shown in Figure 2, we refine the original whole-body pose sequence extracted from video frames (Hu, 2024) into a composite representation, which combines the body pose below the neck with a fixed-size head mask centered on the facial midpoint above the neck. This design circumvents the impact of facial keypoints in original pose sequences on generating synthesized facial movements, thereby enabling lip movements and facial expressions in generated videos to be driven by subsequent audio inputs and other control signals. This composite representation is then mapped through PoseNet and element-wise added to the output of the U-Net’s first convolution layer (Zhang et al., 2024b). In addition, along with extracting image features from the reference image, we detect the facial region and extract an identity embedding using a pre-trained face recognition network (Deng et al., 2018). We draw inspiration from the Cyberhost (Lin et al., 2024) and introduce an additional cross-attention layer after the original cross-attention layer in every U-Net block, specifically designed to generate facial dynamics. Notably, we adopt the strategy from  (Li et al., 2024) to incorporate conditional signal directly related to facial motion, including audio, expression labels, and blink ratio, as an additional control signal that interacts with the identity embedding in the new cross-attention layer. This design mitigates the ambiguity inherent in purely audio-driven approaches, leading to accurate generation of lip motions, facial expressions, and eye blinking while maintaining strong identity consistency.
3.3.2 Training Strategy
@@
We employ a two-stage training strategy to train the rendering module. Given that the original pose-guided SVD backbone (Zhang et al., 2024b) is typically trained on whole-body dance data, we first fine-tune it using collected close-up human talking videos to enhance the generation of magnified facial details. Subsequently, we introduce additional control signals and train the complete network controlled by both audio and pose. In the second stage, we utilize human talking videos with audio-visual synchronization at various scales, including whole-body standing postures, close-up seated postures, and the talking-head scale. For each training videos, we extract audio features as  (Xu et al., 2024a) and randomly select one frame as a reference image to extract image features and identity features (Lin et al., 2024).
4 Experiments
4.1 Setting
@@
Table 1: Quantitative comparison and ablation study of our proposed VersaAnimator on multi-scale animation benchmark, Multi-Animate.
@@
Table 2: Evaluation of text-to-motion capability on the HumanML3D test set, with comparison to state-of-the-art methods such as TM2T (Guo et al., 2022b), T2M (Guo et al., ), MDM (Shafir et al., 2023), MLD (Chen et al., 2023), MotionDiffuse (Zhang et al., 2022b), T2M-GPT (Zhang et al., 2023a), and ReMoDiffuse (Zhang et al., 2023b).
@@
Figure 4: Results on Multi-Animate with different audio and reference images (ranging from head to whole-body).
Metrics. We evaluate the performance of text-to-motion generation using R-Precision (Guo et al., ) and Frechet InceI’m ption Distance (FID) (Heusel et al., 2017), which reflect semantic consistency and the overall motion quality, respectively. For human video generation, we employ a range of metrics. FID, FVD (Unterthiner et al., 2018), SSIM (Wang et al., 2004), and PSNR (Hore & Ziou, 2010) are used to assess low-level visual quality, while E-FID (Deng et al., 2019) is used to evaluate the authenticity of the generated images. CSIM is employed to measure identity consistency. Additionally, we use SyncNet (Prajwal et al., 2020) to calculate Sync-C and Sync-D, which validate the accuracy of audio-lip synchronization.
@@ 4.2 Quantitative Results
Figure 5: Comparison in pose-driven human animation setting. Focus on the background preservation, hand clarity, the occlusion relationship between the hands and the clothing pattern.
Figure 6: Visual illustration of natural head movement.
Comparisons of Detail Preservation. As shown in Figure 5, our method effectively preserves the details of the reference image, such as the pattern on the clothes (green box), the background, and the character’s position. It is worth noting that when the hand movement would be in front of the pattern of the clothes (green box), the part occluded by hands is not shown, in line with the world law.
Comparisons with Pose-Driven Body Methods. Figure 5 also shows that VersaAnimator achieves better clarity in local areas, such as the hands, with additional comparisons provided in the supplementary material.
Visual Illustration of Natural Head Movement. The introduction of 3D human motion allows the person to move as a whole, with the head naturally following the body’s movement, aligning with the natural flow of real human speech. As shown in Figure 5, our method enables natural head movement, enhancing the realism and fluidity.
@@
Figure 7: Qualitative comparison on whether to use code-pose translation module. Left: Stiff and unnatural without pose translation. Right: More natural and fluid motion with pose translation.
4.3 Qualitative Results
@@
Figure 8: Visual illustration of text control for customizing the character’s motion in the generated video.
@@
4.4 Ablation Study
Analysis of Code-pose Translation Module. As shown in Table 1, row 5 indicates the impact of our code-pose translation module. It significantly enhances the quality metric (FID, FVD, SSIM, and PSNR). As shown in Figure 7, this module largely alleviates the stiffness and unreality associated with 3D human modeling.
@@ 5 Conclusion
In this paper, we highlight the challenges faced in the application of human animation, particularly in supporting diverse scenarios that involve different visible areas of human bodies and enabling fine-grained control of body motion by text prompts. We propose VersaAnimator, a versatile audio-driven human video generation method that synthesizes natural and expressive talking human videos, from head to whole-body animation, which allows users to customize the body motion along with well synchronized lip and facial movements. Our comprehensive experiments demonstrate that VersaAnimator outperforms existing methods, enhancing both user-friendly interaction and the overall experience in human animation.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning, with potential applications in Deepfake generation (Mirsky & Lee, 2021).
Appendix A Implementation Details.
For text-controlled, audio-driven motion generation, we set the codebook as 512 × 512, i.e., 512 512-dimension dictionary vectors. Following (Guo et al., 2024; Zhang et al., 2023a), the dataset HumanML3D is extracted into motion features with dimensions 263, which related to local joints position, velocity, and rotations in root space as well as global translation and rotations. The joint number is set to 22. The transformer is composed of 8 transformer layers, with 6 heads and a latent dimension of 384.
We train our diffusion model on 40 hours of self-recorded and web-collected news broadcast videos with synchronized audio and visual content. The dataset is balanced across diverse ethnicities, languages, and shot scales. The average length of the training video clips is 10 seconds. The video diffusion module is trained on 8 NVIDIA A100 GPUs (80GB). The first stage, pose2vid, is trained for 10,000 steps, building upon pre-trained weights from Mimicmotion (Zhang et al., 2024b). The second stage, which incorporates audio input, is trained for an additional 36,000 steps.
Appendix B Comparisons with Pose-driven Body Animation Methods.
Figure 9: Comparisons with pose-driven body animation methods.
Figure 10: Comparisons with pose-driven body animation methods.
We present more comparisons results with pose-driven body animation methods. The visual results in Figure 9 and Figure 10 demonstrate that VersaAnimator maintains superior structural integrity and identity consistency in local regions, such as the hands and face, when compared to the current state-of-the-art methods.
Appendix C Details of Audio Token.
To enhance the encoding of audio for motion-driven animation, we utilized wav2vec (Schneider et al., 2019) as our audio feature encoder. Specifically, we concatenated the audio embeddings from the final 12 layers of the wav2vec model to capture a richer and more diverse range of semantic information across different audio layers. Considering the sequential nature of audio data and its contextual dependencies, we designed a temporal block to extract and fuse the temporal information. Through multiple convolution layers, we transformed the pre-trained audio embeddings into $\left\{c_{\text{audio}}^{t}\right\}_{t=1}^{T/l}$, where $c_{\text{audio }}^{t}\in\mathbb{R}^{D_{a}}$ represents the $t$-th audio token. Here, $l$ is the downsampling rate of the temporal block and ${D_{a}}$ denotes the dimension of the audio token.
Appendix D Comparison of Motion Amplitude.
We observe that the motion generated by other methods exhibits a limited magnitude around the gesture in the reference image. We download several videos from the official VLOGGER website and test our method for comparison. We extract the dwpose frame by frame from the generated video and visualize the dwpose of the upper limbs. The dwpose of all the frames is superimposed on a single image, representing the entire video. As shown in Figure 11, our method covers a larger active area and is more expressive.
Figure 11: Comparison of motion amplitude with VLOGGER (Corona et al., ). We extract the dwpose frame by frame from the generated video and visualize the dwpose of the upper limbs. The dwpose of all the frames is superimposed on a single image, representing the entire video. It can be seen that our method covers a larger active area and is more expressive.
Appendix E Elaborate Version of Related works.
Human Motion Generation. Human motion generation can be broadly categorized into two primary approaches based on the type of input: 1) motion synthesis without conditions (Raab et al., 2022; Tevet et al., 2022; Zhang et al., 2020; Zhao et al., 2020; Yan et al., 2019) and 2) motion synthesis with specified multimodal conditions, such as action labels (Xu et al., 2022; Guo et al., 2020a; Petrovich et al., 2021; Lee et al., 2022; Dou et al., 2023), textual descriptions (Wang et al., ; Chen et al., 2022; Lin & Amer, 2018; Ahn et al., 2018; Tevet et al., ; Zhang et al., 2022a; Tevet et al., 2022; Petrovich et al., 2022; Athanasiou et al., 2022; Lu et al., 2023; Guo et al., ; Ahuja & Morency, 2019), or audio and music (Li et al., b; Tseng et al., 2022; Li et al., a; Dabral et al., 2022; Li et al., 2021b; Siyao et al., ). Due to its user-friendly nature and the convenience of language input, text-to-motion is one of the most important motion generation tasks. Given the remarkable success of diffusion-based generative models in other domains (Rombach et al., 2022), some approaches have employed conditional diffusion models for human motion generation (Zhang et al., 2022b, 2023b, 2024a; Chen et al., 2023). Other works (Zhang et al., 2023a; Guo et al., 2024; Wang, 2023) first discretize motions into tokens using vector quantization (Van Den Oord et al., 2017) and then predict the code sequence of motion.
Appendix F Metric Details and Further Results on Text-to-motion Generation.
Table 3: Evaluation of text-to-motion capability on the HumanML3D test set, with comparison to state-of-the-art methods such as TM2T (Guo et al., 2022b), T2M (Guo et al., ), MDM (Shafir et al., 2023), MLD (Chen et al., 2023), MotionDiffuse (Zhang et al., 2022b), T2M-GPT (Zhang et al., 2023a), and ReMoDiffuse (Zhang et al., 2023b).
[…]
Top 1 Top 2 Top 3
TM2T ${0.424}^{\pm{.003}}$ ${0.618}^{\pm{.003}}$ ${0.729}^{\pm{.002}}$ ${1.501}^{\pm{.017}}$ ${2.424}^{\pm{.093}}$ ${3.467}^{\pm{.011}}$
T2M ${0.455}^{\pm{.003}}$ ${0.636}^{\pm{.003}}$ ${0.736}^{\pm{.002}}$ ${1.087}^{\pm{.021}}$ ${2.219}^{\pm{.074}}$ ${3.347}^{\pm{.008}}$
MDM - - ${0.611}^{\pm{.007}}$ ${0.544}^{\pm{.044}}$ $\mathbf{{2.799}}^{\pm{.072}}$ ${5.566}^{\pm{.027}}$
MLD ${0.481}^{\pm{.003}}$ ${0.673}^{\pm{.003}}$ ${0.772}^{\pm{.002}}$ ${0.473}^{\pm{.013}}$ ${2.413}^{\pm{.079}}$ ${3.196}^{\pm{.010}}$
MotionDiffuse ${0.491}^{\pm{.001}}$ ${0.681}^{\pm{.001}}$ ${0.782}^{\pm{.001}}$ ${0.630}^{\pm{.001}}$ ${1.553}^{\pm{.042}}$ ${3.113}^{\pm{.001}}$
T2M-GPT ${0.492}^{\pm{.003}}$ ${0.679}^{\pm{.002}}$ ${0.775}^{\pm{.002}}$ ${0.141}^{\pm{.005}}$ ${1.831}^{\pm{.048}}$ ${3.121}^{\pm{.009}}$
[…]
Metric details For motion generation, we follow the common metrics of prior works (Guo et al., ) to evaluate the text-to-motion generation performance. Global representations of motion and text descriptions are first extracted with the pre-trained network in (Guo et al., ), and then measured by the following five metrics:
  • R-Precision. Given one motion sequence and 32 text descriptions (1 ground-truth and 31 randomly selected mismatched descriptions), we rank the Euclidean distances between the motion and text embeddings. Top-1, Top-2, and Top-3 accuracy of motion-to-text retrieval are reported.
  • Frechet Inception Distance (FID). We calculate the distribution distance between the generated and real motion using FID (Heusel et al., 2017) on the extracted motion features.
  • Multimodal Distance (MM-Dist). The average Euclidean distances between each text feature and the generated motion feature from this text.
  • Multimodality (MModality). For one text description, we generate 30 motion sequences forming 10 pairs of motion. We extract motion features and compute the average Euclidean distances of the pairs. We finally report the average over all the text descriptions.
We present more metrics for comparison with other methods in Table 3. In addition to the metrics discussed in the main text, we rank second in MM-Dist and MModality, respectively. This demonstrates great generative diversity and multimodal alignment capabilities.
Appendix G Introduction of HumanML3D.
HumanML3D (Guo et al., ) is currently the largest 3D human motion dataset with textual descriptions. The dataset contains 14,616 human motions and 44,970 text descriptions. The entire textual descriptions are composed of 5,371 distinct words. The motion sequences are originally from AMASS (Mahmood et al., 2019) and HumanAct12 (Guo et al., 2020b) but with specific pre-processing: motion is scaled to 20 FPS; those that are longer than 10 seconds are randomly cropped to 10-second ones; they are then re-targeted to a default human skeletal template and properly rotated to face Z+ direction initially. Each motion is paired with at least 3 precise textual descriptions. The average length of descriptions is approximately 12. According to (Guo et al., ), the dataset is split into training, validation, and test sets with proportions of 80%, 5%, and 15%, respectively. We select the best FID model on the validation set and report its performance on the test set.

### v3

@@ 3.3. Generating Photorealistic Talking and Moving Humans with Audio and Motion Control
Figure 4. Results on Multi-Animate with different audio and reference images (ranging from head to whole-body).
[…]
Figure 5. Comparisons of Detail Preservation. Focus on the background preservation, hand clarity, the occlusion relationship between the hands and the clothing pattern.
[…]
Figure 6. Comparisons with pose-driven body animation methods. The red dashed boxes indicate areas with poor performance, such as background changes or unclear hands.
[…]
3.3.1. Co-Speech Human Animation with Video Diffusion Model
@@
We extend an off-the-shelf human animation framework (Zhang et al., 2024a) by incorporating additional conditional signal, including audio, emotion labels, and blinking ratios. This enhancement enables the simultaneous generation of lip movements and whole-body animations.
@@
As shown in Figure 2, we refine the original whole-body pose sequence extracted from video frames (Hu, 2024) into a composite representation, which combines the body pose below the neck with a fixed-size head mask centered on the facial midpoint above the neck. This design circumvents the impact of facial keypoints in original pose sequences on generating synthesized facial movements, thereby enabling lip movements and facial expressions in generated videos to be driven by subsequent audio inputs and other control signals. This composite representation is then mapped through PoseNet and element-wise added to the output of the U-Net’s first convolution layer (Zhang et al., 2024a). In addition, along with extracting image features from the reference image, we detect the facial region and extract an identity embedding using a pre-trained face recognition network (Deng et al., 2019a). We draw inspiration from the Cyberhost (Lin et al., 2024) and introduce an additional cross-attention layer after the original cross-attention layer in every U-Net block, specifically designed to generate facial dynamics. Notably, we adopt the strategy from  (Li et al., 2024b) to incorporate conditional signal directly related to facial motion, including audio, expression labels, and blink ratio, as an additional control signal that interacts with the identity embedding in the new cross-attention layer. This design mitigates the ambiguity inherent in purely audio-driven approaches, leading to accurate generation of lip motions, facial expressions, and eye blinking while maintaining strong identity consistency.
3.3.2. Training Strategy
@@
We employ a two-stage training strategy to train the rendering module. Given that the original pose-guided SVD backbone (Zhang et al., 2024a) is typically trained on whole-body dance data, we first fine-tune it using collected close-up human talking videos to enhance the generation of magnified facial details. Subsequently, we introduce additional control signals and train the complete network controlled by both audio and pose. In the second stage, we utilize human talking videos with audio-visual synchronization at various scales, including whole-body standing postures, close-up seated postures, and the talking-head scale. For each training videos, we extract audio features as  (Xu et al., 2024b) and randomly select one frame as a reference image to extract image features and identity features (Lin et al., 2024).
4. Experiments
4.1. Setting
@@
Table 1. Quantitative comparison and ablation study of our proposed VersaAnimator on multi-scale animation benchmark, Multi-Animate.
@@
Table 2. Evaluation of text-to-motion capability on the HumanML3D test set, with comparison to state-of-the-art methods such as TM2T (Guo et al., 2022c), T2M (Guo et al., 2022b), MDM (Shafir et al., 2023), MLD (Chen et al., 2023), MotionDiffuse (Zhang et al., 2022), T2M-GPT (Zhang et al., 2023c), and ReMoDiffuse (Zhang et al., 2023b).
@@
Table 3. User study results on identity preservation (IP), background preservation (BP), temporal consistency (TC) and visual quality (VQ).
[…]
MimicMotion 30.5% 16.5% 13.5% 33.0%
EchoMimicV2 23.0% 38.0% 47.0% 23.5%
VersaAnimator 96.5% 95.5% 89.5% 93.5%
Metrics. We evaluate the performance of text-to-motion generation using R-Precision (Guo et al., 2022b) and Frechet InceI’m ption Distance (FID) (Heusel et al., 2017), which reflect semantic consistency and the overall motion quality, respectively. For human video generation, we employ a range of metrics. FID, FVD (Unterthiner et al., 2018), SSIM (Wang et al., 2004), and PSNR (Hore and Ziou, 2010) are used to assess low-level visual quality, while E-FID (Deng et al., 2019b) is used to evaluate the authenticity of the generated images. CSIM is employed to measure identity consistency. Additionally, we use SyncNet (Prajwal et al., 2020) to calculate Sync-C and Sync-D, which validate the accuracy of audio-lip synchronization.
@@ 4.2. Quantitative Results
Figure 7. Visual illustration of natural head movement.
[…]
Figure 8. Qualitative comparison on whether to use token2pose translator. Left: Stiff and unnatural without pose translation. Right: More natural and fluid motion with pose translation.
[…]
Figure 9. Visual illustration of text control for customizing the character’s motion in the generated video.
[…]
Comparisons of Detail Preservation. As shown in Figure 5, our method preserves fine details from the reference image, such as clothing patterns, background, and character position. Notably, when hands overlap with clothing (green box), the occluded regions are not rendered, consistent with physical laws.
Comparisons with Pose-Driven Body Methods. Figure 6 also shows that VersaAnimator achieves better clarity in local areas, such as the hands.
Visual Illustration of Natural Head Movement. The integration of 3D motion allows whole-body movement, with the head naturally following, aligning with real human speech. As shown in Figure 7, our method produces natural head motion, enhancing realism and fluidity.
@@
4.3. Qualitative Results
@@
User Study. To evaluate our method against state-of-the-art approaches, we conducted a blind user study with 10 participants. Using 10 reference images and 10 driving videos, we generated 30 clips across three methods. Participants compared video pairs for each input based on visual quality, identity preservation, background preservation, and temporal consistency. Each comparison is repeated $C_{2}^{3}$ times. As shown in Table 3, our method consistently outperformed others across all criteria. More details are provided in the Supplementary Material.
4.4. Ablation Study
Analysis of Token2pose Translator. As shown in Table 1, row 5 indicates the impact of our Token2pose Translator. It significantly enhances the quality metric (FID, FVD, SSIM, and PSNR). As shown in Figure 8, this module largely alleviates the stiffness and unreality associated with 3D human modeling.
@@ 5. Conclusion
In this paper, we highlight key challenges in human animation, particularly the need to support diverse scenarios involving varying portrait scales and the fine-grained control of body motion through text prompts. We introduce VersaAnimator, a versatile talking human framework that generates natural and expressive videos from static portraits. VersaAnimator can animate portraits of various scales while allowing users to customize body movements, all with well-synchronized lip and facial expressions. Extensive experiments demonstrate that VersaAnimator outperforms existing methods, offering more intuitive user interaction and significantly enhancing the overall human animation experience.